1. A feed sample contains 12% moisture. You are feeding a sow 5 lbs per day of this diet. How much. dry matter are you feeding her?
2. A diet you are formulating contains 1.2% calcium. if you are feeding 35 pounds of feed, how much calcium are you feeding in grams?
3. How much salt must you add to a 1 ton batch of feed to provide 3% salt to a finished ration?
Answer #1

1. Moisture % in feed =12%

therefore % of dry matter = 100-12 =88%

then, 5 lbs per day is fed to a cow

Amount of dry matter is 88% of 5lbs

therefore amount of dry matter = ( 88/100)×5 = 4.4lbs

therefore,4.4 lbs dry matter is fed to a cow per day.

2. Simply calcium is 1.2% of 35 pounds = (1.2/100)×35

=0.42 pounds

Now convert 0.42pounds into grams = 0.42 × 453.592

= 190.5 grams

Therefore there is 190.5 grams of calcium in the formulation.

3. We have to find the amount of salt added to 1 ton of feed so that it makes 3% finished ratio.

Let x be the amount of salt in tonnes

3% = (x/1) ×100

x = 3/100

x = 0.03 ton.

Therefore, 0.03 ton of salt should be added to the 1 ton feed to make 3% ratio.